The length of a rectangular pizza is 4 in less than twice its width. The area of the pizza is 160 in^2. Find the dimensions of the pizza.
Answer by mananth(16946)

let the width be x
length = 2x-4
x(2x-4)=160
2x^2-4x-160=0
x^2-2x-80=0
x^2-10x+8x-80=0
x(x-10)+8(x-10)=0
(x-10)(x+8)=0
x=10 inches which is the length
length will be 16 inches